A large tank is placed with water to height H. A small hole is made at base of tank it takes T1 time to decrease the height of water to H/n (n>1); it takes T2 time to take out rest of water. If T1=T2, then the value of n is

Grade:12th pass

2 Answers

Kumari Sanjana
6 years ago
Let A1 be area of cross section of tank and A2 be the area of cross section of hole When water level is x. Then -A1*dx= A2*v*dt-A1.dx=A2.√(2gx).dt(-A1/A2).(dx/√x)=√2g.dtIntegrating both sides-A1/A2.(2/3)((√x)^3=√(2g).tT1= (2A1/3A2.√2g)((√H)^3-(√H/n)^3)T2= (......................)((√H/n)^3)Now ,T1/T2&T1/T2=1 n=(4)^(1/3)
Kumari Sanjana
6 years ago
Let A1 be area of cross section of tank and A2 be the area of cross section of hole, When water level is x. Then -A1*dx= A2*v*dt-A1.dx=A2.√(2gx).dt(-A1/A2).(dx/√x)=√2g.dtIntegrating both sides-A1/A2.(2/3)((√x)^3=√(2g).tT1= (2A1/3A2.√2g)((√H)^3-(√H/n)^3)T2= (......................)((√H/n)^3)Now ,T1/T2&T1/T2=1 n=(4)^(1/3)